### v1

# Analysis of a multi-target linear shrinkage covariance estimator

Abstract
Multi-target linear shrinkage is an extension of the standard single-target linear shrinkage for covariance estimation. We combine several constant matrices - the targets - with the sample covariance matrix. We derive the oracle and a bona fide multi-target linear shrinkage estimator with exact and empirical mean. In both settings, we proved its convergence towards the oracle under Kolmogorov asymptotics. Finally, we show empirically that it outperforms other standard estimators in various situations.
Keywords: Multi-target, linear shrinkage, covariance estimation, general asymptotics
1 Introduction and related work
The covariance matrix plays a major role in numerous machine learning algorithms and statistics. Just to cite a few, array signal processing [1], [2], generalized method of moments estimators [3], financial portfolio management [4], [5] and forecasting [6], functional genomics [7]. However, those algorithms are designed to use the population covariance matrix, which is often unaccessible. The sample covariance matrix severely fails in many applications: for instance, for Markowitz portfolio management it does often not beat a naive uniform diversification among the assets [8].
In the context of Kolmogorov asymptotics, where the ratio of the dimension $p_{n}$ and the number of samples $n$ tends to a finite positive constant $\frac{p_{n}}{n}\rightarrow c>0$, the sample covariance eigenvalue spectrum is spread: high eigenvalues tends at being too high, and low ones, too low [9], [10], [11].
Influenced by the work of Stein on Gaussian mean estimation in 1956 [12], we focus on a specific type of regularization: shrinkage estimation. The motivation is to design a covariance estimator that behaves well in high dimension, with a lower variance than the sample covariance at the cost of an increased bias. This variance-bias trade-off aims at reducing the overall mean-squared error (MSE) of the covariance estimator with respect to the population covariance $\Sigma$. Linear shrinkage estimators linearly combine the sample covariance matrix with a finite set of predefined constant matrices - the targets.
More formally, we denote by $S$ the sample covariance, and we consider $N\in\mathbb{N}$ targets, denoted $(T^{(i)})_{i=1}^{N}$, which are constant matrices of the same size as $S$. Usually, when there is only one target, $N=1$, we chose $T^{(1)}=I$ the identity matrix [13]. With $N=2$ targets, a standard example is the following. Consider the a diagonal block population covariance $\Sigma=\begin{pmatrix}\Sigma_{1}&(0)\\ (0)&\Sigma_{2}\end{pmatrix}.$ Even if $\Sigma_{1}$ and $\Sigma_{2}$ are unknown, we can leverage our knowledge of the block structure, and chose the following targets:
$T^{(1)}=\begin{pmatrix}I_{1}&(0)\\ (0)&(0)\end{pmatrix}\text{ and }T^{(2)}=\begin{pmatrix}(0)&(0)\\ (0)&I_{2}\end{pmatrix}.$ (1)
where $I_{1}$ and $I_{2}$ are the identity matrix of same size as, respectively, $\Sigma_{1}$ and $\Sigma_{2}$. A linear shrinkage covariance estimator $\hat{\Sigma}$ is, by definition:
$\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}.$ (2)
It is a linear combination of the sample covariance and the targets. The challenge is to estimate $(c_{0},...,c_{N})$ only with the observations in order to minimize the Frobenius norm $\lVert\hat{\Sigma}-\Sigma\rVert_{F}^{2}$, where $\Sigma$ is the population covariance.
Many single-target shrinkage estimators (STSE) were proposed, just to name a few: Ledoit-Wolf linear shrinkage [13], its extension for Gaussian distributions named Oracle Approximating Shrinkage (OAS) estimator [14], linear shrinkage with factor models [15], linear shrinkage for elliptical distributions with unknown mean and known radius distribution [16].
However, the performance of those algorithms highly depends on the choice of the target and its similarity to the population covariance $\Sigma$. Thus, there is no general ideal choice of target fitting all covariances $\Sigma$. The use of several targets instead of one can lead to superior classification performance [17, 18]. The methods to compute multi-target linear shrinkage estimators are less numerous. With only two targets, an analytic estimator is given in [17]. The general case for any number of targets an estimator was proposed for Gaussian distributions in [19]. In the distribution-free case, an implicit analytic estimator was proposed in [18], retrieved from solving a strictly convex problem. Nevertheless, to the best of our knowledge, there is no theoretical analysis of any of those algorithms comparable to the quadratic convergences we have in single-target shrinkage.
We address the lack of explicit estimators and theoretical results for multi-target shrinkage covariance estimation. This paper proposes an explicit analytic distribution-free multi-target shrinkage estimator (MTSE). It provides an asymptotic theoretical analysis, proving its quadratic convergence to the oracle estimator in the scope of Kolmogorov asymptotics. Experimentally, our estimator outperforms single-target estimators in the sense of MSE in various situations.
The paper is organized as follows. In section 2, we introduce the notation, definitions and hypotheses necessary to define and address the MTSE problem. In section 3, we expose the problem, the explicit oracle and proposed estimator, and show its quadratic convergence. Section 4 presents the numerical simulations. Finally, section 5 exposes the principal limitations of the approach.
For easier reading, technical proofs of the theoretical results are derived in the appendix, we will only give the idea of proof for intuition in the main corpus.
2 Notation, definitions and hypotheses
Let us introduce the following notation.
Notation 1 (Observations).
We consider a sequence of observation matrices $(X_{n})_{n\in\mathbb{N}^{*}}$ with $X_{n}\in\mathbb{R}^{p_{n}\times n}$ of $n$ iid observations on a system of $p_{n}$ dimensions. Decomposing the covariance matrix, we denote $\Sigma_{n}=\Gamma_{n}\Lambda_{n}\Gamma_{n}^{t}$, where $\Lambda_{n}$ is a diagonal matrix and $\Gamma_{n}$ a rotation matrix. The diagonal elements of $\Lambda_{n}$ are the eigenvalues $\lambda_{1}^{n},...,\lambda_{p_{n}}^{n}$, and the columns of $\Gamma_{n}$ are the eigenvectors $\gamma_{1}^{n},...,\gamma_{p_{n}}^{n}$. $Y_{n}=\Gamma_{n}^{t}X_{n}$ is a $p_{n}\times n$ matrix of $n$ iid observations of $p_{n}$ uncorrelated random variables $(y_{1}^{n},...,y_{n}^{n})$.
Notation 2 (Norm).
[…]
Notation 3 (Convergence).
Let $(E_{n})_{n}$ be a sequence of euclidean spaces with associated norm $\lVert\cdot\rVert_{n}$. The quadratic convergence of a random variable $Z_{n}\in E_{n}$, i.e. $\mathbb{E}[\lVert Z_{n}\rVert_{n}^{2}]\rightarrow 0$, is denoted as $Z_{n}\underset{q.m}{\longrightarrow}0$.
W define the sample covariance in the setting of known and unknown mean, and the set of targets.
Definition 1 (Empirical covariance).
As we are studying both situations of known or unknown mean, we will use a different sample covariance $S_{n}$ depending on the situation. If we know the mean of the distribution, for an observation matrix $X_{n}$ of size $p_{n}\times n$, we define the empirical covariance as:
$S_{n}=(X_{n}-\mathbb{E}[X])(X_{n}-\mathbb{E}[X])^{T}/n.$ (3)
Otherwise, if we do not know the mean of the distribution, we will use the unbiased sample covariance:
$S_{n}=\tilde{X}_{n}\tilde{X}_{n}^{T}/(n-1),\text{ with }(\tilde{X}_{n})_{ik}=(X_{n})_{ik}-\frac{1}{n}\sum_{k^{\prime}=1}^{n}(X_{n})_{ik^{\prime}}.$ (4)
Definition 2 (Set of targets).
For each $n$, we define the finite sequence of $N_{n}\in\mathbb{N}^{*}$ targets: $\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}$, where each $T_{n}^{(k)}$ is a symmetric matrix of size $(p_{n},p_{n})$, not necessarily positive. We assume that for each $n$, $\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}$ form a linearly independent family, orthogonal and normalized.
The “orthogonal and normalized” condition is not really a restriction in practice. Indeed, we can built any linearly independent family $T_{n}^{(k)}$ of symmetric matrices. Then, we can use the Gram-Schmidt algorithm [20] on this family to make it ortho-normalized, without modifying $\text{Span}\left(\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}\right)$, so without changing the space of multi-target estimators of the form $\hat{\Sigma}=c_{0}S+\sum_{i=1}^{N}c_{i}T^{(i)}\text{, where }(c_{0},...,c_{N})\in\mathbb{R}^{N}$.
So, we consider without loss of generality that the family $\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}$ is orthogonal and normalized.
We describe now several assumptions, extensions to our framework of the assumptions used in the linear shrinkage of Ledoit and Wolf [13], which will be used in the following. The first one is the core hypothesis in Kolmogorov asymptotics: the dimension can grow along with the number of samples.
Assumption 1 (Dimension).
There exists a constant $K_{1}$ independent of $n$ such that $p_{n}/n\leq K_{1}$.
Bounded $8$-th moments are necessary for the proofs, even if experimentally we will see that bounded $4$-th moments seem sufficient for the convergence to hold. For the notation, $y_{i1}^{n}$ designates the $i$-th dimension of the first sample in $Y_{n}$. As all samples are i.i.d., it could be any other arbitrary choice of sample in $\llbracket 1,n\rrbracket$ of course. Similarly, $X_{1}^{n}$ is the first sample, of size $p_{n}$, from $X_{n}$.
Assumption 2 (Moments).
There exists a constant $K_{2}$ independent of $n$ such that $\frac{1}{p_{n}}\sum_{i=1}^{p_{n}}\mathbb{E}\left[(y_{i1}^{n})^{8}\right]\leq K_{2}$ and for all $k\in\llbracket 1,N_{n}\rrbracket$, $\mathbb{E}\left[\left\langle X_{1}^{n}(X_{1}^{n})^{T},T_{n}^{(k)}\right\rangle_{n}^{4}\right]\leq K_{2}$.
The number of target $N_{n}$ can depend of $n$, while staying negligible compared to $n$. Intuitively, each target comes with an additional coefficient to estimate, and the estimation error decreases the performance of the bona fide estimator.
Assumption 3 (Number of targets).
The number of targets $N_{n}$ verifies that: $N_{n}=o(n)$.
Already present in [13], this technical assumption, not really restrictive in practice as argued in the same article, is still necessary here.
Assumption 4 (Technical covariance assumption).
With $Q_{n}$ the set of all the quadruples that are made of four distinct integers between $1$ and $p_{n}$, and for all $i\in\llbracket 1,p_{n}\rrbracket$, $\tilde{y}_{i1}^{n}=y_{i1}^{n}-\mathbb{E}[y_{i1}^{n}]$, we have:
$\lim_{n\rightarrow\infty}\frac{p_{n}^{2}}{n}\times\frac{\sum_{(i,j,k,l)\in Q_{n}}(\text{Cov}[\tilde{y}_{i1}^{n}\tilde{y}_{j1}^{n},\tilde{y}_{k1}^{n}\tilde{y}_{l1}^{n}])^{2}}{|Q_{n}|}=0.$
Finally, for the MTSE problem to have a.s. a unique solution, we need the following characterization. In practice, if $\lVert S_{n}\rVert_{n}^{2}-\sum_{k}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}=0$, i.e. if $S_{n}\in\text{Span}\left(\left(T_{n}^{(k)}\right)_{k=1}^{N_{n}}\right)$, simply return $S_{n}$.
Assumption 5 (Well-posed problem).
For each $n\in\mathbb{N}^{*}$, using the associated probability $\mathbb{P}_{n}$ of the distribution indexed by $n$,
$\mathbb{P}_{n}\left[\lVert S_{n}\rVert_{n}^{2}-\sum_{k}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}=0\right]=0.$
3 Oracle and proposed estimator
Remark 1.
As all the estimators in the following are invariant by translation in the case where we do not know the mean, we can suppose $\mathbb{E}[X]=0$ to make the notation lighter. Indeed, we only work with $(\tilde{X}_{n})_{ik}=(X_{n})_{ik}-\frac{1}{n}\sum_{k^{\prime}=1}^{n}(X_{n})_{ik^{\prime}}$, which is not affected by any change of mean.
Given a sample covariance $S_{n}$, we are looking for a minimizer of the 2-norm in the scope of linear combinations of $S_{n}$ and $T_{n}^{(k)}$.
$\operatorname*{minimize}_{\begin{subarray}{c}c\in\mathbb{R}^{N_{n}+1}\end{subarray}}\left\lVert c_{0}S_{n}+\sum_{k=1}^{N_{n}}c_{k}T_{n}^{(k)}-\Sigma_{n}\right\rVert_{n}^{2}.$ (5)
The minimizer $c^{*}\in\mathbb{R}^{N_{n}+1}$, solve the following equation:
$A_{n}c^{*}=b_{n},\text{ with }$ $A_{n}=\begin{pmatrix}\lVert S_{n}\rVert_{n}^{2}&\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&\langle S_{n},T_{n}^{(N_{n})}\rangle\\ \langle S_{n},T_{n}^{(1)}\rangle_{n}&1&&(0)\\ \vdots&&\ddots&\\ \langle S_{n},T_{n}^{(N_{n})}\rangle&(0)&&1\end{pmatrix},$ (6)
$b_{n}=\left[\langle S_{n},\Sigma_{n}\rangle_{n},\langle T_{n}^{(1)},\Sigma_{n}\rangle_{n},...,\langle T_{n}^{(N_{n})},\Sigma_{n}\rangle_{n}\right]^{T}.$
We have:
$\det(A_{n})=\lVert S_{n}\rVert_{n}^{2}-\sum_{k}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2},$ (7)
$\text{com}(A_{n})^{T}=\begin{pmatrix}1&-\langle S_{n},T_{n}^{(1)}\rangle_{n}&\ldots&-\langle S_{n},T_{n}^{(N_{n})}\rangle\\ -\langle S_{n},T_{n}^{(1)}\rangle_{n}&&&\\ \vdots&&(B_{n})&\\ -\langle S_{n},T_{n}^{(N_{n})}\rangle&&&\end{pmatrix},$
$\text{with }B_{n}=\left(\lVert S_{n}\rVert_{n}^{2}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}\right)I_{N_{n}}+QQ^{T},Q=\left[\langle S_{n},T_{n}^{(i)}\rangle_{n}\right]_{i=1,...,N_{n}}^{T}.$
Under Assumption 5, $A_{n}$ is invertible and we have: $A_{n}^{-1}=\frac{1}{\det(A_{n})}\text{com}(A_{n})^{T}$.
Remark 2.
In the particular case where $A_{n}$ is not invertible (not supposing Assumption 5), remove a target $T_{n}^{(i)}$ such as $\langle S_{n},T_{n}^{(i)}\rangle\neq 0$ and the new built $A_{n}$ will be invertible.
Previous derivations lead to the following oracle estimators.
Definition 3 (Oracle estimators).
The oracle estimators solving the minimization problem 5 are, for $i\in\llbracket 1,N_{n}\rrbracket$:
$c_{0}^{*}=\frac{1}{\det(A_{n})}\left(\langle S_{n},\Sigma_{n}\rangle_{n}-\sum_{k=1}^{N_{n}}\langle S_{n},T_{n}^{(k)}\rangle_{n}\langle\Sigma_{n},T_{n}^{(k)}\rangle_{n}\right),c_{i}^{*}=\langle\Sigma_{n},T_{n}^{(i)}\rangle_{n}-c_{0}^{*}\langle S_{n},T_{n}^{(i)}\rangle_{n}.$ (8)
And we define the oracle estimator of covariance $\Sigma_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}^{*}S_{n}+\sum_{k=1}^{N_{n}}c_{k}^{*}T_{n}^{(k)}\right)$, where $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ is the orthogonal projection on the cone of symmetric positive matrices.
The orthogonal projection $\mathcal{P}_{\mathcal{S}_{n}^{+}}$ at the end does not strictly follows the definition of linear shrinkage. However, as $\Sigma_{n}$ is symmetric positive, the projection step decreases the distance $\lVert\Sigma_{n}^{*}-\Sigma_{n}\rVert_{n}$. This step is not necessary for the convergence results to hold, but it is a free improvement we recommend.
Nevertheless, the oracle estimator still depends on $\Sigma_{n}$, and cannot be computed as it is. We propose a bona fide estimator, and show that asymptotically it converges to the oracle one. The strategy is the following, both for known and unknown mean:
  • For $i\in\llbracket 1,N_{n}\rrbracket$, in $c_{i}^{*}$, $\langle\Sigma_{n},T_{n}^{(i)}\rangle_{n}$ is estimated by $\langle S_{n},T_{n}^{(i)}\rangle_{n}$,
  • $\langle S_{n},\Sigma_{n}\rangle_{n}$ is approximated by $\lVert S_{n}\rVert_{n}^{2}-\underbrace{\mathbb{E}\left[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}\right]}_{\mathbb{V}(S_{n})}$. Note that in expectation both objects are equal. $\mathbb{V}(S_{n})$ is estimated by $\hat{\mathbb{V}}(S_{n})$ defined below,
  • Similarly, for each $k\in\llbracket 1,N_{n}\rrbracket$, $\langle S_{n},T_{n}^{(k)}\rangle_{n}\langle\Sigma_{n},T_{n}^{(k)}\rangle_{n}$ is approximated by $\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}-\underbrace{\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{2}\right]}_{\mathbb{V}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)}$. $\mathbb{V}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)$ is estimated by $\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)$, defined below.
Definition 4 (Known mean - Variance estimators).
We suppose the mean to be known. We define:
$\hat{\mathbb{V}}(S_{n})=\frac{1}{n(n-1)}\sum_{k=1}^{n}\left\lVert X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n}\right\rVert_{n}^{2}.$ (9)
Let $i\in\llbracket 1,N_{n}\rrbracket$. We define:
$\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})=\frac{1}{n(n-1)}\sum_{k=1}^{n}\left\langle X_{\cdot k}^{n}\left(X_{\cdot k}^{n}\right)^{T}-S_{n},T_{n}^{(i)}\right\rangle_{n}^{2}.$ (10)
Definition 5 (Unknown mean - Variance estimators).
We suppose the mean to be unknown. We define:
$\bar{b}_{n}^{2}=\frac{1}{n^{2}}\sum_{k=1}^{n}\left\lVert\frac{n}{n-1}\tilde{X}_{\cdot k}^{n}\left(\tilde{X}_{\cdot k}^{n}\right)^{T}-S_{n}\right\rVert_{n}^{2},$ (11)
$\hat{\mathbb{V}}(S_{n})=\frac{(n-1)^{2}}{(n-2)(n-3)}\bar{b}_{n}^{2}-\frac{1}{n(n-2)}\lVert S_{n}\rVert_{n}^{2}-\frac{n-1}{n(n-2)(n-3)}p_{n}\langle S_{n},I_{p_{n}}\rangle_{n}^{2}.$
Let $i\in\llbracket 1,N_{n}\rrbracket$. We define:
$\bar{b}_{T^{(i)},n}^{2}=\frac{p_{n}}{(n-1)^{2}}\sum_{k=1}^{n}\left\langle\tilde{X}_{\cdot k}^{n}\left(\tilde{X}_{\cdot k}^{n}\right)^{T},T_{n}^{(i)}\right\rangle_{n}^{2}-\frac{1}{n}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n},$ (12)
$\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=q_{n}^{(0)}\bar{b}_{T^{(i)},n}^{2}+q_{n}^{(1)}\left\langle S_{n}T_{n}^{(i)},T_{n}^{(i)}S_{n}\right\rangle_{n}-q_{n}^{(2)}\langle S_{n},T_{n}^{(i)}\rangle_{n}^{2},$
$q_{n}^{(0)}=\frac{(n-1)^{2}}{p_{n}(n-2)(n-3)},q_{n}^{(1)}=\frac{n-1}{p_{n}n(n-2)},q_{n}^{(2)}=\frac{n^{2}-2n-1}{n(n-2)(n-3)}.$
Remark 3 (Implementation trick).
$\bar{b}_{T^{(i)},n}^{2}$ can be made fast to compute in Python with Numpy. We use a complex matrix square root of $T_{n}^{(i)}$, denoted $R_{n}^{(i)}$. We have a $O(p_{n}^{3}+np_{n})$-complexity formula:
$\sum_{k=1}^{n}\left\langle\tilde{X}_{\cdot k}^{n}\left(\tilde{X}_{\cdot k}^{n}\right)^{T},T_{n}^{(i)}\right\rangle_{n}^{2}=\frac{1}{p_{n}^{2}}\sum_{k=1}^{n}\left(\sum_{i=1}^{p_{n}}(R_{n}^{(i)}\tilde{X})_{i,k}^{2}\right)^{2}.$ (13)
Lemma 1 (Coordinate estimators).
Under Assumptions 1 to 3, for all $k\in\llbracket 1,N_{n}\rrbracket$, $\langle S_{n},T_{n}^{(k)}\rangle_{n}$ is an unbiased estimator of $\langle\Sigma_{n},T_{n}^{(k)}\rangle_{n}$. Moreover, it exists $M\in\mathbb{R}_{+}$, independent of $n$ such as $\sum_{k=1}^{N_{n}}\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(k)}\rangle_{n}^{4}\right]\leq\frac{M}{n}K_{1}K_{2}$, where $K_{1}$ and $K_{2}$ are defined in Assumptions 1 and 2.
The proof works around the fact that $\left(T_{n}^{(k)}\right)_{k}$ is an orthonormal family to use Assumption 2 then.
Theorem 1 (Variance estimators).
Denoting $\mathbb{V}(S_{n})=\mathbb{E}[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}]$ and $\mathbb{V}\left(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)=\mathbb{E}\left[\langle S_{n}-\Sigma_{n},T_{n}^{(i)}\rangle_{n}^{2}\right]$, under assumptions 1 to 4, in both situations where the mean is known or unknown for all $i\in\llbracket 1,N_{n}\rrbracket$, $\hat{\mathbb{V}}(S_{n})$ and $\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ are unbiased estimators of $\mathbb{V}(S_{n})$ and $\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n})$ respectively. Moreover, $\mathbb{E}\left[\left(\hat{\mathbb{V}}(S_{n})-\mathbb{V}(S_{n})\right)^{2}\right]\longrightarrow 0$ and $\mathbb{E}\left[\left(\sum_{i=1}^{N_{n}}\hat{\mathbb{V}}(\langle S_{n},T_{n}^{(i)}\rangle_{n})-\mathbb{V}(\langle S_{n},T_{n}^{(i)}\rangle_{n}\right)^{2}\right]\longrightarrow 0$.
The proof of this theorem resides in carefully developing for the first part. The convergences are proven by making appear the assumptions 3 and 4 by Cauchy-Schwarz and Jensen inequalities.
Definition 6 (Bona fide estimator).
The bona fide estimators $c$ for the minimization problem 5 are defined, for $i\in\llbracket 1,N_{n}\rrbracket$, as:
$c_{0}$ $=\min\left((\hat{c}_{0})_{+},1\right),\text{ with }\hat{c}_{0}=\frac{1}{\det(A_{n})}\left(\lVert S_{n}\rVert_{n}^{2}-\hat{\mathbb{V}}(S_{n})-\sum_{k=1}^{N_{n}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}^{2}-\hat{\mathbb{V}}\left(\langle S_{n},T_{n}^{(k)}\rangle_{n}\right)\right)\right),$ (14)
$c_{i}$ $=(1-c_{0})\langle S_{n},T_{n}^{(i)}\rangle_{n},$
The MTSE is defined as: $S_{n}^{*}=\mathcal{P}_{\mathcal{S}_{n}^{+}}\left(c_{0}S_{n}+(1-c_{0})\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}\right).$
The truncation of $c_{0}$ between 00 and $1$ theoretically does not change the asymptotic behavior of the estimator. Geometrically, it ensures that $S_{n}^{*}$ is a convex combination between $S_{n}$ and the matrix $\sum_{i=1}^{N_{n}}c_{i}T_{n}^{(i)}$. Empirically, it improves the robustness of the estimator, particularly in small sample size.
Immediately from Theorem 1, we have the following result.
Corollary 1.
Under assumptions 1 to 4, we have $(\hat{c}_{1}-c_{1}^{*})\det(A_{n})\underset{q.m}{\longrightarrow}0$.
Theorem 2 (Loss convergence).
Under assumptions 1 to 5, $S_{n}^{*}$ quadratically converges to the optimal estimator $\Sigma_{n}^{*}$, i.e. $S_{n}^{*}-\Sigma_{n}^{*}\underset{q.m}{\longrightarrow}0$. Consequently, $S_{n}^{*}$ has asymptotically the same loss as $\Sigma_{n}^{*}$, i.e. $\mathbb{E}\left[\left|\left\lVert S_{n}^{*}-\Sigma_{n}\right\rVert_{n}^{2}-\left\lVert\Sigma_{n}^{*}-\Sigma_{n}\right\rVert_{n}^{2}\right|\right]\rightarrow 0$.
The idea of the proof is to prove the convergence on the estimators without the projection step onto the cone of positive matrices, and use the Lipschitz property of projection to conclude. The initial convergence relies on the Lemma A.1 in [13]. The check of the hypotheses for the use of the Lemma comes essentially from the previous results we proved.
4 Experimental results
This section illustrates the practical consequences of the choice of targets on the estimation, the robustness of the estimator to heavy tails, and an application to real non-stationary time series. The experiments were computed on a Lenovo laptop, Intel i5, 32Go Ram. The performance is measured with the Percentage Relative Improvement in Average Loss (PRIAL):
$\text{PRIAL}(\hat{S}_{n})=\frac{\mathbb{E}[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}]-\mathbb{E}[\lVert\hat{S}_{n}-\Sigma_{n}\rVert_{n}^{2}]}{\mathbb{E}[\lVert S_{n}-\Sigma_{n}\rVert_{n}^{2}]}.$ (15)
4.1 Target set impact: Usefulness of a target
We fix the dimension at $p=50$, and $n=25$, and we study the impact of the number of targets $N_{n}$. The population covariance $\Sigma$ is designed to clearly exhibit a choice of useful targets, and compare the role of targets that give information on the population covariance - i.e. that improve significantly the oracle PRIAL-, to those which give less. $\Sigma$ is built as follows: this is a block diagonal matrix composed of 10 block matrices $(B_{i})_{i=1}^{10}$ of shape $(5,5)$. For each $i\in\llbracket 1,10\rrbracket$, $B_{i}$ is drawn from a Wishart distribution with $5$ degrees of freedom and $\sqrt{11-i}I_{5}$ as scale matrix.
$\Sigma=\begin{pmatrix}B_{1}&&(0)\\ &\ddots&\\ (0)&&B_{10}\end{pmatrix}$ (16)
The underlying distribution of $X$ is a multivariate t-distribution with $\nu=9$.
We consider two sets of targets: one set of “aligned” targets $(T_{1})^{(i)}$, and one set of “misaligned” ones $(T_{2})^{(i)}$. The first one is built as following: the first target $T_{1}^{(1)}=I$, to match Ledoit-Wolf STSE [13], and for $i\in\llbracket 2,10\rrbracket$, the target $T_{1}^{(i)}$ is block-diagonal; where $\Sigma$ has the block $B_{i}$, $T_{1}^{(i)}$ has an Identity block; elsewhere $T_{1}^{(i)}$ has null values.
$T_{1}^{(i)}=\begin{pmatrix}(0)&\ldots&(0)\\ \vdots&(\text{block }i:)I&\vdots\\ (0)&\ldots&(0)\end{pmatrix}$ (17)
The set of target is orthogonalized with the Gram-Schmidt algorithm.
For the second set of targets $(T_{2})^{(i)}$, we break the perfect alignment between $T_{1}^{(i)}$ and the block $B_{i}$. For that, we apply a permutation $\sigma$ on both rows and columns of $T_{1}^{(i)}$, where $\sigma$ is defined as:
$\sigma:j\in\llbracket 1,p\rrbracket\longrightarrow(j+2)\text{ mod}p.$ (18)
For the target $i\in\llbracket 1,10\rrbracket$, at index $(k,l)\in\llbracket 1,p\rrbracket^{2}$, we define $(T_{2}^{(i)})_{\sigma(k),\sigma(l)}=(T_{1}^{(i)})_{k,l}$.
We consider the PRIAL of the MTSE (MTSE_1 and MTSE_2 respectively for target set 1 and 2), the oracle ones, and the Ledoit-Wolf STSE (LW), in function of the number $k$ of targets used, between $k=1$, which is equivalent to the STSE, and $k=10$ where the sets of targets $(T_{1}^{(i)})_{i=1}^{k}$ and $(T_{2}^{(i)})_{i=1}^{k}$ are used completely. The results, with $50000$ draws, are shown in Figure 2. The maximum standard deviation is $4\times 10^{-4}$, error bars were removed.
Figure 1: (left) PRIAL of the MTSE, and the oracle, with target sets 1 and 2, and the Ledoit-Wolf STSE as reference, in function of the number of targets in the experiment of target usefulness.
Figure 2: (right) PRIAL difference between the oracle MTSE and Ledoit Wolf linear shrinkage one (MTSE_o - LW) in orange, and estimator MTSE and Ledoit Wolf linear shrinkage one (MTSE - LW) in blue, in function of the number of targets in the setup of the experiment of target set impact.
It confirms the intuition about what is a useful target: the “aligned” set $(T_{1})^{(i)}$ presents a significant better PRIAL than the second misaligned one $(T_{2})^{(i)}$. This comprehension of what makes a target useful is of utter importance to design the target set in a practical use case. The second conclusion is that even if the misaligned targets are not optimal, they still present a better PRIAL than the STSE using only one target.
4.2 Target set impact: Adding useless targets to a good target set
We keep the previous setting for $p$, $n$, $\Sigma$ and the samples $X$ but we change the set of targets.
The first target is the identity matrix $I$, to match with Ledoit-Wolf STSE. For $i\in\llbracket 2,10\rrbracket$, $T^{(i)}$ is designed according to Equation 17. For target $i\geq 11$, the target is drawn from a Wishart distribution with $p$ degrees of freedom and $I_{p}$ as scale matrix. The latter are supposed to be “useless” regarding the structure of $\Sigma$. The set of targets is then orthogonalized with the Gram-Schmidt algorithm.
We ran the experiment for $5000$ times, the maximum standard deviation on any point of the figure is $7\times 10^{-4}$, we did not plot the boxes for readability. We plot the differences of PRIAL between the MTSE and the Ledoit-Wolf STSE, which has only one target. We do the same for the MTSE oracle. While the difference of PRIAL is positive, it means that adding additional targets improved the PRIAL. The result can be seen in figure 2.
Two distinct behaviors emerge from this experiment. The first $10$ useful targets results in a significant improvement of the PRIAL. For the random targets, they can be tagged as useless as they do not improve significantly the oracle PRIAL. The interesting conclusion is that the estimator PRIAL is only slightly negatively affected but stays above the STSE PRIAL. Consequently, adding good targets gives a high reward, and adding useless targets comes at a low price.
Figure 3: (right) Convergence of $\lVert S_{n}^{*}-\Sigma_{n}^{*}\rVert_{n}^{2}$ for different values of $\nu$, where $\nu$ is the degree of freedom of a multivariate t-Student distribution.
4.3 Heavy-tails robustness
In the case of Ledoit-Wolf STSE, Ledoit and Wolf empirically argue that Assumption 2 about finite eighth moments is in practice not necessary, only finite fourth moments would suffice [13]. Thus, we study the convergence of $S_{n}^{*}$ towards $\Sigma_{n}^{*}$ in a setting of heavy tails. For that, we consider multivariate t-Student distributions with $\nu>2$ degrees of freedom, for different values of $\nu$.
We consider $\Sigma$ block diagonal with $5$ blocks of same size drawn independently from Wishart distributions of $p//5$ degrees of freedom, and the set of targets is made of $5$ block diagonal matrices, target $i$ with the Identity as block $i$ and 00 elsewhere. We compare, for $20000$ independent experiments using t-Students with $\nu\in\{20000,8,4,3,2.5\}$ degrees of freedom. $\nu=20000$, verifying all Assumptions, is a reference; $\nu=8$ is at the edge of Assumption 2, $\nu=4$ is at the edge of the empirical conjecture of Ledoit and Wolf, $\nu=3$ is what we found the be the frontier for the convergence to hold, and $\nu=2.5$ is an extreme case.
This experiment, which results are in Figure 3, strengthens the hypothesis of Ledoit and Wolf [13] saying that bounded fourth moments is experimentally enough.
4.4 Application to non-stationary time series
In order to give the practitioner a real example on an application of the MTSE where a “structure” prior is known and can be used, we consider a classical portfolio management problem in finance.
We use daily log-returns of $323$ assets that stayed in the S&P500 between 2010 and 2022  [^1] Data were retrieved from YahooFinance using yfinance package.. We recall that the log-return $r_{t}$ at day $t$ of a price $s_{t}$ is defined as the log-increment: $r_{t}=\log(s_{t})-\log(s_{t-1})$. So we have a $300$-dimensional time series, with one vector per day.
Those $323$ assets, are classified in $11$ industrial sectors, known as the GICS sectors. The general structural idea we want to use is that log-returns of different GICS sectors are less correlated than within the same sector. We expect to have a covariance structure similar to a stochastic block matrix, not as perfect as in Equation 16 due to other mechanisms, but we expect to take advantage of it.
To evaluate the performance of different covariance estimators, we used the General Minimum Variance - GMV - portfolio problem, a standard method to compare covariance estimators. The idea of GMV is to determine at the beginning of each month $T$ a weight vector $w_{T}\in\mathbb{R}^{323}$, as there are $323$ assets here, and invest $w_{i,T}$ on the asset $i$ for one month: this is our portfolio.
At each day $t$ of the month $T$, around $20$ business days, we collect the log-returns of our portfolio $R_{t}=w_{T}^{T}r_{t}$, and at the end of the month we store the empirical variance of our portfolio during the month: $v_{T}=\mathbb{V}[\{R_{t}\}_{t=T}^{T+20}]$. The weight vector $w_{T}$ is designed to minimize this variance $v_{T}$, under constraint that $w_{T}^{T}1=1$. There is closed-form formula, where $\Sigma$ is the covariance matrix of the random vector $r_{t}$: $w_{T}=\frac{\Sigma^{-1}1}{1^{T}\Sigma^{-1}1}$. The pseudo-inverse is used when not invertible.
This problem depends only on the covariance and is perfect to compare different estimators. The experiment goes as following: at the beginning of each month $T$, for each estimator $\hat{\Sigma}$, we estimate it on the last $K$ months, i.e roughly $20K$ data points, and we compute the weight vector $\hat{w}_{T}=\frac{\hat{\Sigma}^{-1}1}{1^{T}\hat{\Sigma}^{-1}1}$ At the end of the month we store the empirical variance $\hat{v}_{T}$. We do this procedure over $135$ months, between 2010 and 2022, and we compare the Cumulative Variance over every months $\sum_{T=1}^{135}v_{T}$, the results are shown in the Table 1. The better the estimator $\hat{\Sigma}$, the lower is the Cumulative Variance.
The number $K$ of months used to fit the covariance estimators varies between $3$ and $15$. The estimators we compare, along with sample covariance $S$, are:
  • MTSE, where the $11$ targets are chosen to be the Identity matrix on the sub-selection of assets belonging to the same GICS sector, and 00 on the other indexes of the matrix,
  • two STSE: LWO, the corrected Ledoit-Wolf estimator in the setting of unknown mean [21], and OAS, the Oracle Approximated Shrinkage estimator [14],
  • two non-linear shrinkage: ANS, the Analytical Nonlinear Shrinkage estimator [22], and QIS, the Quadratic Inverse Shrinkage estimator [23].
The first remark is that increasing the number of samples, from $3$ to $15$ months, does not improve the numerical performance of the estimators we consider. That comes from the non-stationarity of financial time series, one of its stylized facts [24]. It virtually makes the number of samples low.
In this setting, the additional information we plug in MTSE makes the difference in the experiment, as it consistently has a lower Cumulative Variance than the other considered algorithms.
Estimator $3$ months $4$ months $6$ months $9$ months $12$ months $15$ months
MTSE $\mathbf{7.40\times 10^{-3}}$ $\mathbf{7.13\times 10^{-3}}$ $\mathbf{7.30\times 10^{-3}}$ $\mathbf{7.49\times 10^{-3}}$ $\mathbf{8.12\times 10^{-3}}$ $\mathbf{8.03\times 10^{-3}}$
LWO $7.60\times 10^{-3}$ $7.34\times 10^{-3}$ $7.51\times 10^{-3}$ $7.70\times 10^{-3}$ $8.57\times 10^{-3}$ $8.27\times 10^{-3}$
OAS $7.63\times 10^{-3}$ $7.42\times 10^{-3}$ $7.74\times 10^{-3}$ $8.24\times 10^{-3}$ $9.63\times 10^{-3}$ $9.52\times 10^{-3}$
ANS $7.62\times 10^{-3}$ $7.41\times 10^{-3}$ $7.56\times 10^{-3}$ $7.91\times 10^{-3}$ $1.15\times 10^{-2}$ $2.14\times 10^{-2}$
QIS $3.73\times 10^{-2}$ $7.83\times 10^{-3}$ $8.56\times 10^{-3}$ $1.04\times 10^{-2}$ $1.68\times 10^{-2}$ $4.90\times 10^{-2}$
S $5.55\times 10^{-2}$ $7.28\times 10^{-2}$ $5.84\times 10^{-2}$ $6.89\times 10^{-2}$ $8.95\times 10^{-2}$ $4.04\times 10^{-1}$
Table 1: GMV Cumulative Variance for different estimators and number of months to fit them.
5 Limitations
The strongest assumptions made in the theoretical analysis are that we have $n$ i.i.d. samples and bounded eighth moments. In practice, the i.i.d. assumption is hard to assess, particularly in time series analysis, as we saw in the financial dataset. Eighth moments can be a limitations, even if we saw that experimentally the convergence seems to hold for lower moments.
As in single-target estimation, the central point for performance in a task is a meticulous choice of targets, and general methods to design them still lack in the literature to the best of our knowledge.
6 Conclusion
We derived the oracle estimator solving the multi-target linear shrinkage covariance estimator and proposed an explicit analytic estimator of it. The main result of the contribution is the proof of the convergence of the estimator under Kolmogorov asymptotics, while the number of targets remains negligible compared to the number of samples.
The numerical simulations highlight the gain in performance to add useful targets to the estimation. Experimentally, we argue that the estiamtor convergence holds with only bounded fourth moments. The experiment in a non-stationary setup underlines the potential of adaptability of multi-target estimation in those situations, compared to other linear or non-linear shrinkage methods.